OpenStudy (brooke..help00):

I will medal and fan. This is more of a personal question so I can understand this better. So I am working on linear equations and graphing them. Can anyone please list some examples of something that I can substitute y=mx+b with. Please also give an explanation. Thanks so much🙂

OpenStudy (johnweldon1993):

Hmm not sure what you mean? If I give you a point on a line and the slope of the line, could you tell me the equation of the line? Similarly...If I give you the equation of a line, would you be able to tell me if a certain point is on that line? And if I gave you the 'x' and 'y' intercepts of a line, would you be able to tell me the slope of that line and then the equation of the line?

OpenStudy (brooke..help00):

Can you give me an equation and explain how to get the y intercept and slope?

OpenStudy (blazeryder):

An easy one: \[-20x+5y=10\] First get the x variable on the left side of the equation. To do that add 20x to both sides. \[-20x+20x+5y=20x+10\]\[5y=20x+10\]Now you need to get the y by itself. To do that we divide by 5. \[\frac{ 5y }{ 5 }=\frac{ 20x }{ 5 } + \frac{ 10 }{ 5 }\]\[y=4x+2\] So, the slope is 4 and the y-intercept is 2 I hope this helps and I also hope this is what you were looking for. =)

OpenStudy (brooke..help00):

Yes! Thank you so much! I am so happy you explained:)

OpenStudy (blazeryder):

No problem! Glad I could help!